# Theories Without Models: Uncontrolled Idealizations in Particle Physics

### v1

@@ Abstract.
The perturbative treatment of realistic quantum field theories, such as quantum electrodynamics, requires the use of mathematical idealizations in the approximation series for scattering amplitudes. Such mathematical idealisations are necessary to derive empirically relevant models from the theory. Mathematical idealizations can be either controlled or uncontrolled, depending on whether current scientific knowledge can explain whether the effects of the idealization are negligible or not. Drawing upon negative formal results in asymptotic analysis (failure of Borel summability) and renormalization group theory (failure of asymptotic safety), we argue that the mathematical idealizations applied in perturbative quantum electrodynamics should be understood as uncontrolled. This, in turn, leads to the problematic conclusion that such theories do not have theoretical models in the standard understanding of this term. The existence of unquestionable empirically successful theories without theoretical models has significant implications both for our understanding of the theory-model relationship in physics and the concept of empirical adequacy.
@@ 1. Theories Without Models: An Argument
    Perturbative quantum field theory is best understood as a framework for theories where individual theories are picked out within the framework via particular action principles.
@@
    For a theory to have theoretical models is for a representation of the quantities predicted by the theory to be derivable either: i) as deductive consequences of the theory; or ii) as a result of an approximate derivation based upon controlled mathematical idealizations.
@@
    Contemporary particle physics features perturbative quantum field theories, such as quantum electrodynamics, in which the derivation of predictions requires an ineliminable appeal to uncontrolled idealizations.
@@
    Contemporary particle physics includes theories, such as quantum electrodynamics, without theoretical models.
P1 we take to be uncontroversial and straightforwardly justified by contemporary scientific practice. P2 is a standard definition drawn from the philosophical literature and we take it to be also uncontroversial, although upon to contestation and revision with good reason. P3 is the controversial premise, and much of the paper will be devoted to defending it in the context of the well-known formal challenges to the rigorous formulation of perturbative quantum field theories. Given our arguments in favour of P3 are accepted, and one wishes to reject the conclusion, either on the grounds of being unintuitive or due to other negative implications, the obvious implication via contraposition is then that we should revise our definition of theoretical models to reject P2.
@@
Before we proceed to situate our analysis in the various background literatures, let us make the aims of this paper completely clear by indicating what this paper is not about. Here we are not arguing that particle physics does not have models per se, nor that there is a fundamental or basic inconsistency or inadequacy in the framework of perturbative quantum field theory. Indeed, models of various non-theoretical sorts abound in particle physics and there exist various important formal results that indicate precisely in which circumstances the idealizations involved in perturbative quantum field theory can be understood to be controllable. Our aim here is to, first, demonstrate that in the context of unquestionably empirically successful examples of individual perturbative quantum field theories applied in particle physics, such as quantum electrodynamics, the conditions for the controllability of the idealizations in question can be expected to fail and to, second, point out that this implies that these theories do not have models in the standard sense of theoretical model. We take this conclusion to be the result of the combination of uncontroversial results in perturbative quantum field theory and standard definitions in the philosophy of science and, as such, to be all the more worth of interest given the modesty of the individual argumentative steps involved.
@@ 2.1. Mathematical Idealizations in Particle Physics
Mathematical idealizations are formal simplifications of the structure of a mathematical model or system of equations that aim to facilitate the provision of explicit analytical or numerical solutions. The classic simple example of a mathematical idealization is the small angle approximation in the solution of the differential equation for pendulum motion and we will return to this example later. In modern physics, mathematical idealizations are often based upon truncation series expansion. For example, approximating the exponential function by the first few terms of the Taylor expansion, i.e. $e^{x}\approx 1+x+\frac{x^{2}}{2}$. Mathematical idealizations should be distinguished from other forms of idealizations, much discussed in contemporary philosophy, which typically focus on the processes of constructing model systems or representations by distorting or omitting features of their target systems. [^1] In his seminal paper on idealization McMullin (1985) explicitly distinguishes between ‘mathematical idealization’ and ‘construct idealization’. The discussions in Pincock (2007) and Batterman (2009) mainly concern the former type in which the focus is on the mathematical treatment of the models, while discussions in Weisberg (2007), Morrison (2015) and Portides (2021) mainly concern the latter. See also Norton (2012) for related discussion.
@@
A controlled idealization is an idealization for which all the necessary resources are available in current scientific knowledge to explain why the effect of the introduced assumptions in a model is, up to required degree, negligible. [^2] Our deployment of this concept builds upon the idea of a ‘controllable idealization’ due to (Sklar 2000, pp.44-5). See also Batterman (2005, 2014); Wayne (2011); King (2016); Knuuttila and Morgan (2019). A controlled mathematical idealization is then a simplification of the formal structure of a mathematical model or system of equations in which all the necessary resources are available in current scientific knowledge to explain why the effects of the relevant formal simplifying assumption are, up to a required degree, negligible. Typically, in controlling a mathematical idealization a scientist is able to provide estimates of the deviation between the theoretical predictions of the idealized model and the exact solutions of the original mathematically intractable theoretical model. In what follows, we will distinguish between the situation where we do not at present have the resources to control an idealization but we have good reason to expect that this is only a practical limitation and the situation where we have good reason to believe that we will never be in the situation to control the idealization in question. The first we will call uncontrolled idealizations and the second uncontrollable idealizations.
@@
The standard model of particle physics is a marvel of empirical science and has been variously tested to astounding degrees of accuracy. The framework within which the standard model is presented is perturbative Quantum Field Theory (QFT). [^3] There have been various attempts to cast the standard model in terms of non-perturbative quantum field theory. However, to date, such attempts have been met with limited success. Here we focus exlcusively on perturbative approaches with the framework of Lagrangian Quantum Field Theories and neglect analysis of models in Algebraic Quantum Field Theory (AQFT) approaches. An issue of particular interest, worthy of consideration in future work, is the implications for our analysis of work on perturbative approaches to AQFT (Rejzner 2016). Following Wallace (2006, 2018) and Koberinski (2021) our analysis is based on an understanding of perturbative QFT as a general theoretical framework of principles and general mathematical constraints from which certain well-defined Lagrangian quantum field theories describing particular interactions are constructed, such as quantum electrodynamics (QED) and quantum chromodynamics (QCD). [^4] Here we deviate slightly from the terminology of Koberinski (2021), since he prefers to refer to the Lagrangian QFTs as dynamical models, although he admits that he uses the term ‘dynamical model’ in the way most would use the term ‘theory’. These Lagrangian QFTs necessarily come with an implicit cut-off limit implying that they are effective field theories describing phenomena within a particular range of energy scales. The connection of these theories with the empirical data is achieved via the construction of empirically relevant models of specific interactions within the theory’s domain (e.g., a meson-meson scattering for QED), usually by perturbation methods. In what follows, quantum electrodynamics is thus referred to as the ‘background theory’ from which empirically relevant models of certain interactions are derived with the use of approximation techniques.
@@
The perturbative treatment of the class of realistic models in quantum field theories requires an approximation of the scattering amplitudes in terms of a divergent power series of the coupling parameter. The divergence occurs (i) from the fact that the total sum of the series is not known to converge or known not to converge and/or (ii) from the fact that each individual term of the series diverges at very high and very low energies leading to the so-called ultraviolet and infrared divergences respectively. As will be shown, the necessary mathematical treatment of these challenges in the case of important examples of QFTs, such as quantum electrodynamics, requires the use of uncontrolled idealisations which break the formal link needed to identify the background theory with its theoretical models. In particular, the justification for the truncation of the series after the first few terms can only be established if the series is strongly asymptotic, by showing that it is Borel summable, cf. Miller (2023). Correspondingly, the justification for the renormalizability of a quantum field theory via renormalization group methods can only be achieved by showing that a theory is asymptotically safe by calculating the corresponding beta-function, cf. Huggett (2002). However, these requirements have not been demonstrated to hold in important examples of QFTs, and there is evidence that they do not in fact hold in QED at least. Thus, the introduced mathematical idealizations are uncontrolled, and plausibly they should be understood as uncontrollable.
@@
The general understanding of perturbative quantum field theory that we adopt is close in spirit to recent discussions on the role of theories, models, approximations and renormalization techniques in quantum field theories. We conclude this subsection by giving a brief overview of this recent related literature. In an analysis related to but distinct from our own, Fraser (2020) argues that perturbative methods in quantum field theory do not produce models, rather they only produce approximations of physical quantities like cross-sections. In more recent work, Fraser’s view on the role of approximations in quantum field theory has been further elaborated by de Olano et al. (2022) by showing via certain historical examples that approximations are instrumental in determining the empirical content of perturbative quantum field theories and should be recognised as a distinct and equally important category of theoretical output compared to idealized systems. de Olano et al. (2022) further stress that the integral role of approximations in assigning empirical and physical content to a model cannot be easily accounted for by extant philosophical approaches to scientific modelling. Koberinski (2021) also provides a historical study of the mathematical developments of Yang-Mills theories to highlight the fact that quantum field theories are rarely axiomatic systems with a neat set of deductive consequences, but are rather modular and rely heavily on key conceptual and mathematical tools that can be treated independently. Similarly, Miller (2023) argues that the state-space semantics developed by Beth (1960) and Fraassen (1970) fails to capture perturbatively evaluated observables in cases of divergent series, and proposes a slight modification that allows the accommodation of these results. These results complement and support our own approach. As such, we take ourselves to be operating in the mainstream of the philosophy of quantum field theory, despite the counter-intuitiveness of our conclusion.
@@ 2.2. Models and Theories
Under what might be called the ‘canonical view’ on theoretical models, to be a model of a theory is to stand in the relation to the theory that is equivalent to that between a logical model and a formal theory expressed as a set of axioms or formal sentences. Such a view of theoretical models is common to all major approaches to the structure of scientific theories. That is, whether or not one holds the semantic view of theories under which theories are identified with the theoretical models, it is still standard to consider the models of a theory as an important aspect of a theory’s structure and to define such theoretical models via the canonical view. [^5] See, for example, Savage et al. (1990); French and Ladyman (1999); Lutz (2012); Winther (2021). Most vividly, the models of general relativity are unambiguously specified by parings of stress-energy tensors and spacetime geometries that solve the Einstein Field Equations. A highly influential articulation of the canonical view is due to van Fraassen (1980), who indicates that theoretical models should be understood as deductive products of the background theory in the sense that they are derived from a set of fundamental equations: ‘[t]he sense in which a theory offers or presents us with a family of models – the theoretical models – is just the sense in which a set of equations presents us with the set of its own solutions […] When the equations formulate a scientific theory, their solutions are the models of that theory’ (ibid. p.311).
@@
We thus end up with a deeply unintuitive situation where we have theories without models. On close inspection the situation turns out to be not just unintuitive but also concerning since theoretical models can be expected to play a number of important roles within scientific theories. A feature of particular significance is that on most accounts it is the theoretical models that allow theories to make contact with empirical evidence. This feature is not only pivotal to van Fraassen’s constructive empiricist view of science but can be expected to be central to any account of how scientific theories are applied in practice. In particular, the empirical adequacy of a theory is standardly understood to assessed by the predictive success of the theoretical models it provides. Thus we arrive at the significant worry that the existence of theories without models not only undermines the pristine picture of theories and theoretical models, but also indicates that the empirical adequacy of theories such as quantum electrodynamics cannot be cogently expressed. We will return to this worry in Section 6.
@@
The autonomy of models from theories as a criticism to the semantic view of theories has of course been discussed in the past in other contexts. [^6] For instance, Portides (2005) criticises the semantic view by showing how the liquid drop model of the nucleus was introduced in the 1930s to explain – amongst other things – the Weizsäcker semi-empirical formula of the binding energy of the nucleus. Portides convinsingly argued that certain crucial assumptions of the model lack the necessary theoretical justifications which would allow one to consider the liquid drop model as a theoretical model (ibid., p.1294). In a similar spirit, Cartwright et al. (1995) and Morrison (1999) argued for the autonomy of models from theories using the examples of the Londons’s model of superconductivity and Prandtl’s model of ideal fluids respectively. However, such arguments typically appeal to examples of phenomenological models in physics. The main characteristic of phenomenological models is that their construction is primarily guided by empirical observation and experimental data, maintaining minimal dependence on the background theoretical framework. As a result, such models often violate basic theoretical principles and fundamental laws of the background physical theory, and hence, phenomenological models exhibit a type of non-theoretical models in virtue of their independence and autonomy from the theory. Our worry is rather different. That is, our claim is that there is a very different type of non-theoretical models in physics – for which the models of perturbative quantum field theory are a prime example – in which the relationship between theories and models breaks down not because these models are primarily guided by phenomenological considerations; but rather, because the mathematical treatment of these models results to the introduction of uncontrollable idealizations due to the lack of a rigorous mathematical or physical justification.
@@
The class of problematic models that we identify is similar to what Redhead (1980) calls floating models. For Redhead, these are models that are disconnected from a background theory due to a computational gap arising from the scientist’s inability to justify the validity of the mathematical approximations used for their construction, but at the same time, unlike phenomenological models, fail to provide successful empirical predictions. They are thus, in a sense, ‘floating’ in that they are detached both from theory and the experimental data. Their value is to be found in the fact that they often serve as preliminary steps for probing various essential features of the theory and exploring the possible ways of refining it in order to yield further empirically successful models. However, unlike Redhead’s floating models, the models of perturbative quantum field theory exhibit remarkable agreement with experimental data despite the fact that they are not connected to the background theory via a secure approximate derivation based upon a controlled mathematical idealizations. This is, we contend, a deeply challenging situation that has not been adequately addressed in the vast literature on theories and models in science.

### v2

@@ Abstract.
The perturbative treatment of realistic quantum field theories, such as quantum electrodynamics, requires the use of mathematical idealizations in the approximation series for scattering amplitudes. Such mathematical idealizations are necessary to derive empirically relevant models from the theory. Mathematical idealizations can be either controlled or uncontrolled, depending on whether current scientific knowledge can explain whether the effects of the idealization are negligible or not. Drawing upon negative mathematical results in asymptotic analysis (failure of Borel summability) and renormalization group theory (failure of asymptotic safety), we argue that the mathematical idealizations applied in perturbative quantum electrodynamics should be understood as uncontrolled. This, in turn, leads to the problematic conclusion that such theories do not have theoretical models in the natural understanding of this term. The existence of unquestionable empirically successful theories without theoretical models has significant implications both for our understanding of the theory-model relationship in physics and the concept of empirical adequacy.
@@ 1. Theories Without Models: An Argument
    Perturbative quantum field theory is best understood as a framework for theories where individual theories are picked out within the framework via (for example) their field content, symmetries, and actions.
@@
    For a theory to have theoretical models is for a representation of the quantities predicted by the theory to be derivable either: i) as deductive consequences of the theory; or ii) as a result of an approximate derivation based upon controlled mathematical idealizations within the theory.
@@
    Contemporary particle physics features perturbative quantum field theories, such as quantum electrodynamics, in which we have good reasons to believe the derivation of predictions requires an ineliminable appeal to uncontrolled idealizations.
@@
    Contemporary particle physics includes theories, such as quantum electrodynamics, which we have good reasons to believe are theories without theoretical models.
P1 we take to be uncontroversial and straightforwardly justified by contemporary scientific practice in which quantum electrodynamics and quantum chromodynamics are treated as distinct theories within the framework of quantum field theory and not as models of quantum field theory. P2 is the natural definition of models of a theory in physical context and is consistent with that used in the philosophical literature (van Fraassen 1980; French and Ladyman 1999; Winther 2021). We take it to be also uncontroversial, although open to contestation and revision with good reason. The term ‘theoretical model’ is used here as a term of art and its precise meaning will be clarified below. P3 is the controversial premise, and much of the paper will be devoted to defending it in the context of the well-known formal challenges to the rigorous formulation of perturbative quantum field theories. Given our arguments in favour of P3 are accepted, and one wishes to reject the conclusion, either on the grounds of being unintuitive or due to other negative implications, the obvious implication via contraposition is then that we should revise our definition of theoretical models to reject P2.
@@
Before we proceed to situate our analysis in the various background literatures, let us make the aims of this paper completely clear by indicating what this paper is not about. Here we are not arguing that particle physics does not have models per se, nor that there is a fundamental or basic inconsistency or inadequacy in the framework of perturbative quantum field theory. Indeed, models of various non-theoretical sorts abound in particle physics and there exist various important mathematical results that indicate precisely in which circumstances the idealizations involved in perturbative quantum field theory can be understood to be controllable.
This article serves two main goals. The first is to demonstrate that in the context of unquestionably empirically successful examples of individual perturbative quantum field theories applied in particle physics, such as quantum electrodynamics, the conditions for the controllability of the idealizations in question can be expected to fail. The second is to point out that this implies that these theories do not have models in the straightforward sense of theoretical model. We take this conclusion to be the result of the combination of uncontroversial results in perturbative quantum field theory and straightforward definitions in the philosophy of science and, as such, to be all the more worthy of interest given the modesty of the individual argumentative steps involved.
@@ 2.1. Mathematical Idealizations in Particle Physics
Mathematical idealizations are formal simplifications of the structure of a mathematical model or system of equations that aim to facilitate the provision of explicit analytical or numerical solutions. The classic simple example of a mathematical idealization is the small angle approximation in the solution of the differential equation for pendulum motion and we will return to this example later. In modern physics, mathematical idealizations are often based upon the truncation of series expansions. For example, approximating the exponential function by the first few terms of the Taylor expansion, i.e. $e^{x}\approx 1+x+\frac{x^{2}}{2}$. Mathematical idealizations should be distinguished from other forms of idealizations, much discussed in contemporary philosophy, which typically focus on the processes of constructing model systems or representations by distorting or omitting features of their target systems. [^1] In his seminal paper on idealization McMullin (1985) explicitly distinguishes between ‘mathematical idealization’ and ‘construct idealization’. The discussions in Pincock (2007) and Batterman (2009) mainly concern the former type in which the focus is on the mathematical treatment of the models, while discussions in Weisberg (2007), Morrison (2015) and Portides (2021) mainly concern the latter. See also Norton (2012) for related discussion.
@@
A controlled idealization is an idealization for which all the necessary resources are available in current scientific knowledge to explain why the effect of the introduced assumptions in a model is, up to required degree, negligible. [^2] Our deployment of this concept builds upon the idea of a ‘controllable idealization’ due to Sklar (2000, pp.44-5). See also Batterman (2005, 2014); Wayne (2011); King (2016); Knuuttila and Morgan (2019). A controlled mathematical idealization is then a simplification of the formal structure of a mathematical model or system of equations in which all the necessary resources are available in current scientific knowledge to explain why the effects of the relevant formal simplifying assumption are, up to a required degree, negligible. Such resources may come either from within the theoretical tools of a given background physical theory in which the idealizations occur, or, when this is not possible, by appeal to different physical theories in which case one has what we shall call an exogenous justification.
Examples of mathematical idealizations in physics that, when all goes well, are controlled include linearity assumptions, infinite differentiability assumptions, and adiabaticity assumptions (i.e. one function is assumed to change slowly with respect to another). Such idealizations are controlled when scientists are able to provide reasons why the effects of the relevant formal simplifying assumption are, up to a required degree, negligible. Typically, in controlling a mathematical idealization a scientist is able to provide estimates of the deviation between the theoretical predictions of the idealized model and the exact solutions of the original mathematically intractable theoretical model.
In what follows, we will distinguish between the situation where we do not at present have the resources to control an idealization but we have good reason to expect that this is only a practical limitation and the situation where we have good reason to believe that we will never be in the situation to control the idealization in question. The first we will call uncontrolled idealizations and the second uncontrollable idealizations.
@@
The standard model of particle physics is a marvel of empirical science and has been variously tested to astounding degrees of accuracy. The framework within which the standard model is presented is perturbative Quantum Field Theory (QFT). [^3] There have been various attempts to cast the standard model in terms of non-perturbative quantum field theory. However, to date, such attempts have been met with limited success. Here we focus exclusively on perturbative approaches with the framework of Lagrangian Quantum Field Theories and neglect analysis of models in Algebraic Quantum Field Theory (AQFT) approaches. For the classical philosophical discussion of the relative merits of perturbative and algebraic approaches see Wallace (2006); Fraser (2011, 2009); Wallace (2011). An issue of particular interest, worthy of consideration in future work, is the implications for our analysis of work on perturbative approaches to AQFT (Rejzner 2016). Following Wallace (2006, 2018) and Koberinski (2021) our analysis is based on an understanding of perturbative QFT as a general theoretical framework of principles and general mathematical constraints from which certain well-defined Lagrangian quantum field theories describing particular interactions are constructed, such as quantum electrodynamics (QED) and quantum chromodynamics (QCD). [^4] Here we deviate slightly from the terminology of Koberinski (2021), since he prefers to refer to the Lagrangian QFTs as dynamical models, although he admits that he uses the term ‘dynamical model’ in the way most would use the term ‘theory’. These Lagrangian QFTs necessarily come with an implicit cut-off limit implying that they are effective field theories describing phenomena within a particular range of energy scales. The connection of these theories with the empirical data is achieved via the construction of empirically relevant models of specific interactions within the theory’s domain (e.g., a meson-meson scattering for QED), usually by perturbation methods. In what follows, quantum electrodynamics is thus referred to as the ‘background theory’ from which empirically relevant models of certain interactions are derived with the use of approximation techniques.
@@
The perturbative treatment of the class of realistic models in quantum field theories requires an approximation of the scattering amplitudes in terms of a divergent power series of the coupling parameter. The divergence occurs (i) from the fact that the total sum of the series is not known to converge or known not to converge and/or (ii) from the fact that each individual term of the series diverges at very high and very low energies leading to the so-called ultraviolet and infrared divergences respectively. As will be shown, the necessary mathematical treatment of these challenges in the case of important examples of QFTs, such as quantum electrodynamics, requires the use of uncontrolled idealizations which break the formal link needed to identify the background theory with its theoretical models. In particular, in cases where the corresponding functions are unknown like in perturbative QED, the justification for the truncation of a series expansion after the first few terms can be established if the series is strongly asymptotic, by showing that it is Borel summable, cf. Miller (2023). Correspondingly, the justification for the renormalizability of a quantum field theory via renormalization group methods can be achieved by showing that a theory is asymptotically safe by calculating the corresponding beta-function, cf. Huggett (2002). However, these requirements have not been demonstrated to hold in important examples of QFTs, and there is evidence that they do not in fact hold in QED at least. Thus, the introduced mathematical idealizations are uncontrolled, and plausibly they should be understood as uncontrollable.
It is important to note, in this context that the formal challenges regarding the establishing of the renormalizability of a quantum field theory are not ameliorated by appeal to the Effective Field Theory (EFT) framework and the formal control such a framework offers over perturbatively non-renormalizable terms. Indeed, the ‘Wilsonian’ renormalization group approach, which we will consider in detail later, is one of the principal formal foundations of the EFT paradigm. Moreover, the other key formal foundation, and the basis of the crucial property of ‘decoupling’ between effective and high-energy theories, is the Appelquist-Carazzone theorem (Appelquist and Carazzone 1975) and this assumes an underlying renormalizable theory (with different mass scales) (Hartmann 2001). Furthermore, following the recent arguments of Franklin (2020), it is (effective) renormalizability that accounts for the effectiveness of effective field theories. The EFT framework is an illustration of rather than an alternative to the need for justification for renormalizability in QFT.
Given the inability to control the idealizations within the theory, a possible route towards the justification of the mathematical idealizations of QED is to look to physics beyond the standard perturbative analysis and particle content (e.g. introduce ‘instanton’ effects to remedy the failure of Borel summability) or beyond the standard model entirely (e.g. consider implications from asymptotically safe quantum gravity theories to deal with the failure of asymptotic safety). While this is a fruitful way of controlling idealizations in physics, when such exogenous justifications appeal to different physical theories, they cannot establish that the resulting models are, strictly speaking, models of the theory of QED. As will be shown in the next subsection, under the natural sense of ‘models of a theory’ relevant to physics, a theoretical model is a model that can be deductively derived from the background theory modulo some controllable idealizations within the theory – i.e. using resources from the tools of the theory plus mathematics and logic. QED, as it currently stands, is unable to provide its own models due to the inability of controlling the idealizations involved. Most significantly, even if the establishment of a UV completion within a quantum gravity scenario ultimately turns out to be successful, the new resulting models will ultimately not be models of QED, but rather models of the conjunction of QED and the new theory of quantum gravity, given that all the introduced idealizations are controllable. [^5] This feature points to a crucial disanalogy between the notion of a ‘model of a theory’ as understood in the context of physical theories and as applied in model theory. Whereas, in model theory if $M$ is a model of theory $T+T^{\prime}$ it is also a model of theory $T$, such situations do not in general obtain in physical theory. Most explicitly, we could consider $T$ to be special relativity and $T^{\prime}$ to be general relativity. There are then clearly models of $T+T^{\prime}$ which are not models of $T$. Thanks to an anonymous reviewer for pointing this issue out to us. This is orthogonal to our conclusions, however. As it stands, and when understood as a theory of the interactions of charged particles with the electromagnetic field, QED does not provide theoretical models in the natural sense of ‘models of a theory’ relevant to physics. We shall return to this issue in Section 5.
The general understanding of perturbative quantum field theory that we adopt is close in spirit to earlier discussions on the role of theories, models, approximations and renormalization techniques in quantum field theories. We conclude this subsection by giving a brief overview of this literature.
To our knowledge, the first philosophical discussion of the specific problem of justifying the deductive relations between theories and models in QFT can be found in the wide-ranging and hugely insightful article of Hartmann (2001). Based upon a detailed discussion of QCD and EFTs the author notes that ‘there might not be a ‘controlled’ deductive relation between the model and an underlying theory, and if a derivation of the model from the theory is actually carried through, further assumptions have to be made to obtain the model, and these assumptions (which might turn out to be more dubious than the assumptions made by the original model) again require a justification, and so on, ad infinitum’ (p. 293). We take this analysis to be very much in line with our own appraisal of the theory-model relation for the case of QED.
@@
In a more recent analysis, related to but distinct from our own, James Fraser argues that perturbative methods in quantum field theory do not produce models, rather they only produce approximations of certain physical quantities such as cross-sections and decay widths (Fraser 2020) . In more recent work, Fraser’s view on the role of approximations in quantum field theory has been further elaborated together with co-authors in de Olano, Fraser, Gaudenzi, and Blum (2022) by showing, via certain historical examples, that approximations are instrumental in determining the empirical content of perturbative quantum field theories and should be recognised as a distinct and equally important category of theoretical output compared to idealized systems. de Olano, Fraser, Gaudenzi, and Blum (2022) further stress that the integral role of approximations in assigning empirical and physical content to a model cannot be easily accounted for by extant philosophical approaches to scientific modelling.
Koberinski (2021) also provides a historical study of the mathematical developments of Yang-Mills theories to highlight the fact that quantum field theories are rarely axiomatic systems with a neat set of deductive consequences, but are rather modular and rely heavily on key conceptual and mathematical tools that can be treated independently. Similarly, Miller (2023) argues that the state-space semantics developed by Beth (1960) and Fraassen (1970) fails to capture perturbatively evaluated observables in cases of divergent series, and proposes a slight modification that allows the accommodation of these results. These results complement and support our own approach. As such, we take ourselves to be operating in the mainstream of the philosophy of quantum field theory, despite the counter-intuitiveness of our conclusion.
@@ 2.2. Models and Theories
Under what might be called the ‘canonical view’ on theoretical models, to be a model of a theory is to stand in the relation to the theory that is analogous to that between a logical model and a formal theory expressed as a set of axioms or formal sentences. Such a view of theoretical models is common to most major approaches to the structure of scientific theories. That is, whether or not one holds the semantic view of theories under which theories are identified with the theoretical models, it is still standard to consider the models of a theory as an important aspect of a theory’s structure and to define such theoretical models via the canonical view. [^6] See, for example, Savage et al. (1990); French and Ladyman (1999); Lutz (2012); Winther (2021). The natural sense of ‘models of a theory’ that is relevant to physics draws on an analogy to the use of model in model theory but is a clear and independent conceptualisation. [^7] We note that some approaches to the semantic view invest the analogy with deeper significance. The most vividly and logically rigours example is the models of general relativity are unambiguously specified by pairings of stress-energy tensors and Lorentzian four metrics that solve the Einstein Field Equations. A highly influential articulation of the canonical view is due to van Fraassen (1980), who indicates that theoretical models should be understood as deductive products of the background theory in the sense that they are derived from a set of fundamental equations: ‘[t]he sense in which a theory offers or presents us with a family of models – the theoretical models – is just the sense in which a set of equations presents us with the set of its own solutions […] When the equations formulate a scientific theory, their solutions are the models of that theory’ (ibid. p.311).
@@
We thus end up with a deeply unintuitive situation where we have theories without models. On close inspection the situation turns out to be not just unintuitive but also concerning since theoretical models can be expected to play a number of important roles within scientific theories. A feature of particular significance is that on most accounts it is the theoretical models that allow theories to make contact with empirical evidence. This feature is not only pivotal to van Fraassen’s constructive empiricist view of science but can be expected to be central to any account of how scientific theories are applied in practice. In particular, the empirical adequacy of a theory is standardly understood to be assessed by the predictive success of the theoretical models it provides. Thus, we arrive at the significant worry that the existence of theories without models not only undermines the pristine picture of theories and theoretical models, but also indicates that the empirical adequacy of theories such as quantum electrodynamics cannot be cogently expressed. We will return to this worry in Section 6.
@@
The autonomy of models from theories as a criticism to the semantic view of theories has of course been discussed in the past in other contexts. [^8] For instance, Portides (2005) criticises the semantic view by showing how the liquid drop model of the nucleus was introduced in the 1930s to explain – amongst other things – the Weizsäcker semi-empirical formula of the binding energy of the nucleus. Portides convinsingly argued that certain crucial assumptions of the model lack the necessary theoretical justifications which would allow one to consider the liquid drop model as a theoretical model (ibid., p.1294). In a similar spirit, Cartwright et al. (1995) and Morrison (1999) argued for the autonomy of models from theories using the examples of the Londons’s model of superconductivity and Prandtl’s model of ideal fluids respectively. However, such arguments typically appeal to examples of phenomenological models in physics. The main characteristic of phenomenological models is that their construction is primarily guided by empirical observation and experimental data, maintaining minimal dependence on the background theoretical framework. As a result, such models often violate basic theoretical principles and fundamental laws of the background physical theory. Phenomenological models are thus a type of non-theoretical models in that they are independent and autonomous from the theory. Our worry is rather different. That is, our claim is that there is a very different type of non-theoretical models in physics – for which the models of perturbative quantum field theory are a prime example – in which the relationship between theories and models breaks down not because these models are primarily guided by phenomenological considerations; but rather, because the mathematical treatment of these models requires the introduction of uncontrollable idealizations due to the lack of a rigorous mathematical or physical justification.
@@
The class of problematic models that we identify is similar to what Redhead (1980) calls floating models. For Redhead, these are models that are disconnected from a background theory due to a computational gap arising from the scientist’s inability to justify the validity of the mathematical approximations used for their construction, but at the same time, unlike phenomenological models, fail to provide successful empirical predictions. They are thus, in a sense, ‘floating’ in that they are detached both from theory and the experimental data. Their value is to be found in the fact that they often serve as preliminary steps for probing various essential features of the theory and exploring the possible ways of refining it in order to yield further empirically successful models. However, unlike Redhead’s floating models, the models of perturbative quantum field theory exhibit remarkable agreement with experimental data despite the fact that they are not connected to the background theory via a secure approximate derivation based upon a controlled mathematical idealization. This is, we contend, a deeply challenging situation that has not been adequately addressed in the vast literature on theories and models in science.